# 1st Dimitrios Bouzoulas Energy and Mechanical Engineering

### v1

@@ V Discussion
In this paper, an automated labelling method for low-light pedestrian detection leveraging corresponding IR and RGB images was presented. Different baseline YOLOv11 models were tested for both the IR image-based autolabelling and label evaluation through the training of RGB detection model with the autolabels. Training with the autolabels yielded higher mAP@50 and mAP@50-95 than training with the ground truth labels in 6 out of 9 cases.
@@
Figure 4: KAIST ground truth edge case labels, where pedestrians are poorly visible in the RGB frame.
The proposed autolabel performance exceeding ground truth labels in some cases can be attributed to edge cases in ground truth labels where pedestrians are extremely poorly visible in the RGB images due to occlusion and poor illumination (Fig. 4). Such samples can cause the model to overfit to the training data, as generalizable detection features are not found.
@@
The proposed approach was only verified using the KAIST dataset and further research is needed. The next steps include the gathering of a larger IR-RGB dataset to evaluate the performance consistency in larger datasets as well as the effects of a larger number of edge case scenarios with occlusion and poor illumination. Additionally, state-of-the-art methods in infrared detection should be tested as the proposed pipeline relies solely on infrared detections and their quality.
As the proposed method’s results showed potential for low-light applications, similar approaches can be applied to other low-light tasks. One such task is pose estimation, as it could be implemented in a similar manner to the proposed method. Such research could aid the development of more complex systems such as pedestrian crossing intention predictions enabling more advanced decision-making in low-light conditions. Another promising future application is detecting animals, like deer, in low-light conditions.

### v2

@@ V Discussion
In this paper, an automated labelling method for low-light pedestrian detection leveraging corresponding IR and RGB images was presented. YOLOv11l, RF-DETR, and Faster RCNN architectures were tested for both the IR image-based autolabelling and lowlight RGB detection, trained with the autolabeled data. Training with the autolabels yielded higher mAP@50 and LAMR than training with the ground truth labels in 5 out of 6 cases and higher mAP@50-95 in all cases.
@@
Figure 5: KAIST ground truth difficult case labels, where pedestrians are poorly visible in the RGB frame.
Training with the autolabels yielded better results than training with the ground truth labels as shown in Table III and demonstrated on the testing data in Figure 4. This difference in performance could be attributed to the hard cases in ground truth labels, where pedestrians are extremely poorly visible in the RGB images due to occlusion and poor illumination, with examples of such cases shown in Figure 5. Such samples can cause the model to overfit to the training data, as generalizable detection features are not found, resulting in lower overall performance. Another possible explanation is the difficulty of the low-light RGB detection task. The mAP scores of the IR models are up to two times higher compared to the RGB models. Likely, the accuracy of the autolabels is not the main bottleneck, but instead the limitations of RGB detection in low-light conditions.
In our experiments, the model architecture was found to play a major role in the final performance. All IR models were trained on the same data, DETR outperforming YOLO and RCNN with a significant margin (Table II). A similar pattern was observed when low-light RGB models were trained on the ground truth, DETR being the best-performing model (Table III). These results underline the importance of model selection, as DETR was able to adapt to both infrared and low-light conditions significantly better using the same data as the other models.
Another parameter explored in this paper was the autolabelling detection confidence threshold. When the confidence threshold was lowered from 0.5 to 0.25, Recall increases and Precision decreases for all models, while the F1 score increases for IR-YOLO and IR-RCNN and maintains a similar value for IR-DETR. (Table II). The low-light RGB models trained on the autolabels generated with a confidence threshold of 0.25 performed the best in all cases (Table III). There seemed to be a correlation between the F1 score of the autolabels and the performance of the resulting RGB prediction model.
Overall, all of the low-light RGB models have a low recall rate, with the highest being 0.369 for DETR trained on the IR-DETR labels with 0.25 confidence. This highlights both the need for more quality training data and the potential upside from sensor fusion approaches. Camera-based detection is highly affected by lighting conditions, while less lighting-sensitive sensors can provide reliable additional sources of information. Finally, this approach does not pose computation time limitations, allowing for the usage of ensemble models for autolabelling.
The proposed approach was only verified using the KAIST dataset, and further research is needed. The next steps include the gathering of a larger IR-RGB dataset to evaluate the performance consistency in larger datasets, as well as the effects of a larger number of edge case scenarios with occlusion and poor illumination. Additionally, state-of-the-art methods in infrared detection should be tested as the proposed pipeline relies solely on infrared detections and their quality.
@@
As the proposed method’s results showed potential for low-light applications, similar approaches can be applied to other low-light tasks. One such task is pose estimation, as it could be implemented in a similar manner to the proposed method. Such research could aid the development of more complex systems, such as pedestrian crossing intention prediction, enabling more advanced decision-making in low-light conditions.